OpenStudy (anonymous):

Which of the following combinations of reactants will result in a synthesis reaction? (2 points) Select one: a. Cl2 + HBr b. Ca + H2O c. MgO + H2O d. C2H2 + O2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

c MgO + H2O = Mg(OH)2

OpenStudy (anonymous):

are you sure its C

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

cl2 + 2HBR= 2HCL+ BR2 Ca + 2H2O = ca(oh)2 + H2 2C2H2+5 O2= 4CO2+ 2H2O